# Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis

### v1

@@ Abstract
The upcoming era of large-scale, high-cadence astronomical surveys demands efficient and robust methods for time-series analysis. ARIMA models provide a versatile parametric description of stochastic variability in this context. However, their practical use is limited by the challenge of selecting optimal model orders while avoiding overfitting. We present a novel solution to this problem using a Bayesian framework for time-series modelling in astronomy by combining Autoregressive Integrated Moving Average (ARIMA) models with the Nested Sampling algorithm. Our method yields Bayesian evidences for model comparison and also incorporates an intrinsic Occam’s penalty for unnecessary model complexity. A vectorized ARIMA–Nested Sampling framework is implemented allowing us to perform model selection across grids of Autoregressive (AR) and Moving Average (MA) orders, with efficient inference of selected model parameters. The method is validated on simulated and real astronomical time series, including the yearly sunspots number record, Kepler lightcurve data of the red giant KIC 12008916, and TESS photometry of the exoplanet host star Ross 176. In all cases, the algorithm correctly identified the true or best-fitting model while simultaneously yielding well-constrained posterior distributions for the model parameters. Our results demonstrate that Nested Sampling offers a potentially rigorous alternative to autoregressive model selection in astronomical time-series analysis.
@@
methods: statistical, methods: data analysis, (Sun:) sunspots, exoplanets.
pubyear: 2025pagerange: Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis6
@@ 1 Introduction
Time series analysis is a crucial tool in the field of time domain astronomy. As next generation astronomical surveys and telescopes are set to yield unprecedented volumes of complex time series data, an efficient preliminary method for their analysis is necessary. Some examples of such methods are frequency domain analysis (Lomb, 1976; Scargle, 1982), Gaussian Processes (Foreman-Mackey et al., 2017) and Machine Learning methods (Richards et al., 2011; Baron, 2019). Within this diverse landscape of methodologies, parametric autoregressive modelling offers a complementary approach to time series analysis in the form of ARIMA models (Elorrieta, Felipe et al., 2019; Akhter et al., 2020; Carruba & Aljbaae, 2021).
@@
The ARIMA (Autoregressive-Integrated-Moving Average) framework is a suite of models employed in analysing and forecasting time series data from various domains including, but not limited to, economics, finance and climate science. The Autoregressive (AR) component of ARIMA was first introduced by Yule (1927) to study the number of sunspots. These models capture the autocorrelation in a time series by linearly regressing the present value over its own past (lagged) values. Moving Average (MA) models on the other hand, operate on a similar principle by expressing the present value as a linear combination of past forecast errors or residuals. The idea of systematically combining these modelling methods can be traced back to Box & Jenkins (1976), who formalized them into ARIMA with the introduction of the Integrated (I) component to handle non-stationary time series. These are labelled as ARIMA(p,d,q)(p,d,q) models, characterized by the respective orders $p$, $d$ and $q$ of the Autoregressive (AR), Integrated (I) and Moving Average (MA) components.
@@
ARIMA models are seldom used in the analysis of astronomical time series data. The primary reason being that they require evenly sampled data in time, which is often in contrast to ground based astronomical datasets due to their irregular cadences. Nevertheless, ARIMA modelling can still be applied, with reasonable success, by binning light curve data with moderately irregular cadences (Feigelson et al., 2018). Upcoming facilities such as the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), the Roman Space Telescope (Spergel et al., 2015; Johnson et al., 2020), and ongoing space missions like TESS (Ricker et al., 2015) and Gaia (Gaia Collaboration et al., 2016) will also produce adequately regular and high cadence time series data suitable for ARIMA modelling . The generality and flexibility of the ARIMA framework allows it to efficiently model a diverse range of astronomical time series. However, this very flexibility comes with the risk of over-parametrization and overfitting. It is non-trivial to choose the optimal $(p,d,q)$ order of the ARIMA model. The heuristic diagnostics approach taken by the standard Box-Jenkins methodology, used for ARIMA model specification and validation, may limit robustness in some cases. Model selection on the basis of the Akaike Information Criterion (AIC) (Akaike, 1974) and Bayesian Information Criterion (BIC) (Schwarz, 1978) rely on maximum likelihood optimization, which may bias the model selection, especially for complex likelihood shapes.
@@
The Nested Sampling algorithm used in Bayesian computation has shown promise in model selection problems in the context of Astrophysics (Trotta, 2008). Recent advances in GPU accelerated Nested Sampling (Yallup et al., 2025) have reduced the computational costs associated to this method, allowing it to be implemented for a variety of model selection problems (Ormondroyd et al., 2025; Lovick et al., 2025; Leeney et al., 2025; Prathaban et al., 2025a; Yallup, 2025). Using Nested Sampling in conjunction with ARIMA models can thus provide an alternative solution to the problem of selecting the right $(p,d,q)$ order, with the added benefit of also returning posterior samples for the ARIMA parameters. The previously discussed risk of overfitting is avoided with an in-built Occam’s penalty imparted on the Bayesian evidences of over-parametrized model fits.
@@
In the following section, we discuss the mathematical formalism of ARIMA models and review the nested sampling algorithm. Section 3 then describes the methods and details of using the Nested Sampling algorithm for ARIMA model selection. The results of applying this framework on artificial and real astronomical time series data are presented in Section 4. We discuss potential future directions for this work and end with conclusions in Section 5.
@@ 2 Background
Subsection 2.1 outlines the mathematical background and properties of ARIMA models. For a more comprehensive exposition of ARIMA models, readers are directed to the standard texts of Box & Jenkins (1976). In subsection 2.2, we discuss the nested sampling algorithm and its role in model selection.
@@ 2.1 ARIMA Framework
An ARMA process fundamentally assumes the time series data to be stationary. A time series is defined to be stationary if its statistical properties, such as mean, variance and autocorrelation, remain constant over time. The Integrated (I) component of ARIMA models is an added utility to deal with non-stationary time series using finite differencing. Its associated order $d$ is the number of times the raw observations are differenced to render the series stationary. After fitting the differenced time series to an ARMA process, the forecast values $\hat{y}_{t}$ are integrated back to recover the original sequence trend.
@@ 2.1.1 Stationarity and Invertibility Constraints
Even though the problem of non-stationarity in the time series data $y_{t}$ is solved by differencing, the ARIMA model fitted to the data must itself produce a stationary and stable time series forecast. This requirement manifests in the form of two mathematical constraints on the AR and MA weights $\phi_{a}$ and $\theta_{m}$ - the stationarity and invertibility constraints, respectively. Mathematically, these constraints are defined by imposing the following condition: all roots of the AR and MA characteristic polynomials - $\Phi^{p}(\mathcal{B})$ and $\Theta^{q}(\mathcal{B})$, must lie outside the unit circle. Therefore an ARIMA(p,d,q)(p,d,q) process is stationary and invertible if the absolute value of all the roots of $\Phi^{p}(\mathcal{B})$ and $\Theta^{q}(\mathcal{B})$ are greater than one.
The stationarity constraint on the AR weights ensures that the modelled time series exhibits a stable and mean-reverting behaviour. It allows for the influence of past values to eventually decay with time preventing the series from diverging to infinity. This can be demonstrated for a simple AR(1) model, for which the characteristic polynomial is
@@
$\epsilon_{t}=y_{t}-\theta_{1}y_{t-1}-\theta_{2}^{2}y_{t-2}-\theta_{3}^{3}y_{t-3}\ldots$ (15)
@@ 2.2 Nested Sampling
The Nested Sampling algorithm, first developed by Skilling (2006), is primarily used to compute the evidence term in Bayesian inference. Given a dataset $D$, and a model $\mathcal{M}$ characterized by parameters $\theta_{i}$, Bayesian inference updates the prior distribution on the model parameters $P(\theta_{i}|\mathcal{M})$ to obtain the posterior distribution $P(\theta_{i}|D;\mathcal{M})$ using the Bayes theorem:
@@
It is generally intractable to analytically evaluate this integral, especially in higher dimensional spaces. Nested Sampling is a numerical algorithm which transforms this multi-dimensional evidence integral into a one dimensional problem by introducing the prior volume - the amount of fractional prior mass contained within an iso-likelihood contour of $\mathcal{L^{\prime}}$ in the prior space:
@@ 2.2.1 Sampler Output and Details
$\mathcal{D}_{\mathrm{KL}}$ is the Kullback-Leibler divergence (Kullback & Leibler, 1951) which represents the amount of concentration of the posterior relative to the prior:
@@
$n_{\text{live}}$ is the number of live points sampled from the prior. The total convergence time of the nested sampling run scales directly with $\mathcal{D}_{\mathrm{KL}}$ and $n_{\text{live}}$. For a fixed uncertainty $\sigma_{\log{Z}}$, $n_{\text{live}}$ is directly proportional to $\mathcal{D}_{\mathrm{KL}}$. The number of live points $n_{\text{live}}$ is therefore chosen to optimize precision and time for convergence.
@@ 2.2.2 Model Selection
The evidence $Z_{i}$ associated with any model is thus a principled measure of the goodness of fit. This in turn is related to the likelihood function $\mathcal{L}(D|\theta_{i})$ as well as the size of the prior space over which this fit is carried. The latter factor ultimately results in the Occam’s razor that penalizes the evidence of an overfitted model.
@@ 3 Methodology
This section outlines the methodology adopted to apply the nested sampling algorithm on ARIMA models. We discuss details of the likelihood and prior functions in subsection 3.1, the nested sampler in subsection 3.2, and finally the model selection and validation procedures in subsections 3.3 and 3.4, respectively.
The following quantities are treated as the set of parameters characterizing any given ARIMA model : the respective AR and MA weights $\phi_{a}$ and $\theta_{m}$, the missing initial state ($D_{0}$ and $\epsilon_{0}$), the standard deviation $\sigma$ associated to the present error $\epsilon_{t}$ and the unconditional mean of the time series $\mu$, which can be shown to be related to the intercept term $c$ through:
@@ A Gaussian likelihood function is chosen for the $n$ observations of the given time series $D_{t}$ :
After fitting the ARIMA model, the resulting residuals are assumed to be independent Gaussian random variables with mean zero and constant variance ($\sigma^{2}$). Since the noise model enters explicitly through the likelihood function, the ARIMA forecast values $\mathbf{\hat{y}_{t}}$, used in computing the likelihood should be completely deterministic. These deterministic forecasts $\mathbf{\hat{y}_{t}}$ are obtained from Equation 3 by setting $\epsilon_{t}=0$. The missing initial values for the time series data $D_{0}$ and the forecast errors $\epsilon_{0}$ are treated as nuisance parameters in our inference.
@@
We now discuss the choice of prior distributions for the ARIMA model parameters. For $\sigma$, a reasonable choice of prior is a Half-Normal or a Half-Cauchy distribution with scale $\sigma^{\prime}$. We adopt a relatively weak prior for $\sigma$ by choosing a half-normal distribution truncated from below at zero and scaled by $\sigma^{\prime}=50$. The prior choice for the unconditional mean $\mu$ depends on the time series data under consideration. For a differenced or de-trended time-series, a standard normal distribution is chosen as prior for $\mu$. In other cases, a wide normal distribution centred around the expected long-term mean $\mu_{0}$ of the time-series is used. The prior distribution for the $p$ missing initial values of $D_{0}$ is set equivalent to the prior for $\mu$, as we expect these values to lie close to $\mu$ for a stationary series. For the initial residuals $\epsilon_{0}$, a standard normal distribution is chosen. In the case of an ARIMA($1,d,1$) model, a uniform distribution bounded between -1 and 1 is the most straightforward choice of priors for the coefficients $\phi_{a}$ and $\theta_{m}$. The bounds are in accordance with the stationarity and invertibility conditions discussed in subsection 2.1.1. However, for higher order ARIMA models, the constraints on the individual weights are coupled and hence the choice of prior distributions for them is non-trivial.
@@
For this paper, we assign a mathematically constrained normal prior distribution $\mathcal{N}(0,\sigma^{\prime\prime})$ on the ARMA weights and implement a “rejection sampling" approach. The scale of the distribution $\sigma^{\prime\prime}$ is normally set to one as we do not expect the time-series under consideration to be very strongly auto-correlated. An initial set of $10^{3}\times n_{\text{live}}$ points is sampled from this distribution. The points are tested for stationarity and invertibility by calculating the roots of the AR and MA characteristic polynomials, and imposing the conditions discussed in subsection 2.1.1. The parameter points passing this test are then added to a pool of valid particles. This process is continued iteratively until the desired number of valid live particles $n_{\text{live}}$ is reached in this pool. The normal prior distributions on the weights are also constrained to the stationary and invertible regions of the parameter space.

### v2

@@ Abstract
The upcoming era of large-scale, high-cadence astronomical surveys demands efficient and robust methods for time-series analysis. ARIMA models provide a versatile parametric description of stochastic variability in this context. However, their practical use is limited by the challenge of selecting optimal model orders while avoiding overfitting. We present a novel solution to this problem using a Bayesian framework for time-series modelling in astronomy by combining Autoregressive Integrated Moving Average (ARIMA) models with the Nested Sampling algorithm. Our method yields Bayesian evidences for model comparison and also incorporates an intrinsic Occam’s penalty for unnecessary model complexity. A vectorized ARIMA–Nested Sampling framework with GPU-acceleration support is implemented, allowing us to perform model selection across grids of Autoregressive (AR) and Moving Average (MA) orders, with efficient inference of selected model parameters. We validate the approach using simulated time series with known ground-truth parameters and demonstrate accurate recovery of both model order and parameters. We then apply the method to several astronomical datasets, including the historical sunspot number record, stellar light curves of KIC 12008916 and Kepler 17 from the Kepler mission, and quasar light curves of 3C 273 and S4 0954+65 from the TESS mission. In all cases, the ARIMA models selected by this method were able to accurately model the stochastic variability in the time series data. Our results demonstrate that nested sampling offers a rigorous and computationally tractable alternative to autoregressive model selection in astronomical time-series analysis.
@@
methods: statistical, methods: data analysis, (Sun:) sunspots, stars: activity, exoplanets, (galaxies:) quasars: general
pubyear: 2026pagerange: Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis14
@@ 1 Introduction
Time series analysis is a crucial tool in the field of time domain astronomy. As next-generation astronomical surveys and telescopes are set to yield unprecedented volumes of complex time series data, a robust and scalable method for their analysis is necessary. Some examples of such methods are frequency-domain analysis (Lomb, 1976; Scargle, 1982), Gaussian Processes (Foreman-Mackey et al., 2017) and Machine Learning methods (Richards et al., 2011; Baron, 2019). Within this diverse landscape of methodologies, parametric autoregressive modelling offers a complementary approach to time series analysis in the form of ARIMA models (Elorrieta, Felipe et al., 2019; Akhter et al., 2020; Carruba & Aljbaae, 2021).
@@
The ARIMA (Autoregressive-Integrated-Moving Average) framework is a suite of models employed in analysing and forecasting time series data from various domains including, but not limited to, economics, finance and climate science. The Autoregressive (AR) component of ARIMA was first introduced by Yule (1927) to study the number of sunspots. These models capture the autocorrelation in a time series by linearly regressing the present value on its own past (lagged) values. Moving Average (MA) models on the other hand, operate on a similar principle by expressing the present value as a linear combination of past forecast errors or residuals. The idea of systematically combining these modelling methods can be traced back to Box & Jenkins (1976), who formalized them into ARIMA with the introduction of the Integrated (I) component to handle non-stationary time series. These are labelled as ARIMA(p,d,q)(p,d,q) models, characterized by the respective orders $p$, $d$ and $q$ of the Autoregressive (AR), Integrated (I) and Moving Average (MA) components.
@@
ARIMA models are seldom used in the analysis of astronomical time series data. The primary reason is that they require evenly sampled data in time, which is often in contrast to ground-based astronomical datasets due to their irregular cadences. Nevertheless, ARIMA modelling can still be applied, with reasonable success, by binning light curve data with moderately irregular cadences (Feigelson et al., 2018). Upcoming facilities such as the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), the Roman Space Telescope (Spergel et al., 2015; Johnson et al., 2020), and ongoing space missions like TESS (Ricker et al., 2015) and Gaia (Gaia Collaboration et al., 2016) will also produce adequately regular and high-cadence time series data suitable for ARIMA modelling. The generality and flexibility of the ARIMA framework allows it to efficiently model a diverse range of astronomical time series. However, this very flexibility comes with the risk of over-parametrization and overfitting. It is non-trivial to choose the optimal $(p,d,q)$ order of the ARIMA model. The heuristic diagnostic approach taken by the standard Box-Jenkins methodology, commonly used for ARIMA model specification and validation, may limit robustness in some cases. Model selection on the basis of the Akaike Information Criterion (AIC) (Akaike, 1974) and Bayesian Information Criterion (BIC) (Schwarz, 1978) relies on maximum likelihood optimization, which may bias the model selection, especially for complex likelihood shapes.
@@
The Nested Sampling algorithm used in Bayesian computation has shown promise in model selection problems in the context of astrophysics (Trotta, 2008). Recent advances in GPU-accelerated nested sampling (Yallup et al., 2025) have reduced the computational costs associated with this method, allowing it to be implemented for a variety of model selection problems (Ormondroyd et al., 2025; Lovick et al., 2025; Leeney et al., 2025; Prathaban et al., 2025a; Yallup, 2025). While the use of Bayesian inference for autoregressive models has been explored in the econometrics literature (Barnett et al., 1996; Marriott et al., 1996), such approaches have relied on MCMC sampling methods. Using nested sampling in conjunction with ARIMA models can thus provide an alternative solution to the problem of selecting the optimal $(p,d,q)$ order, with the added benefit of also returning posterior samples for the ARIMA parameters. The previously discussed risk of overfitting is avoided with an in-built Occam’s penalty imposed on the Bayesian evidences of over-parametrized model fits. The main contribution of our work is a vectorized ARIMA-Nested Sampling framework with GPU-acceleration support which can be applied to diverse astronomical time series datasets.
@@
In the following section, we present a review of the mathematical formalism of ARIMA models and the nested sampling algorithm. Section 3 then describes the methodology adopted, details of using the nested sampling algorithm for ARIMA model selection, and defines the criterion for a successful model fit. The results of applying this framework on artificial and real astronomical time series data are presented in Section 4. We discuss potential future directions for this work and end with conclusions in Section 5.
@@ 2 Background
Subsection 2.1 outlines the mathematical background and properties of ARIMA models. For a more comprehensive exposition of ARIMA models, readers are directed to the standard text of Box & Jenkins (1976). In subsection 2.2, we discuss the nested sampling algorithm and its role in model selection.
@@ 2.1 ARIMA Framework
An ARMA process fundamentally assumes the time series data to be stationary. A time series is defined to be stationary if its statistical properties, such as mean, variance and autocorrelation, remain constant over time. The Integrated (I) component of ARIMA models is an additional tool for dealing with non-stationary time series using finite differencing. Its associated order $d$ is the number of times the raw observations are differenced to render the series stationary. After fitting the differenced time series to an ARMA process, the forecast values $\hat{y}_{t}$ are integrated back to recover the original sequence trend.
@@ 2.1.1 Stationarity and Invertibility Constraints
Even though the problem of non-stationarity in the time series data $y_{t}$ is solved by differencing, the ARIMA model fitted to the data must itself produce a stationary and stable time series forecast. This requirement manifests in the form of two mathematical constraints on the AR and MA weights $\phi_{a}$ and $\theta_{m}$ - the stationarity and invertibility constraints, respectively. Mathematically, these constraints are defined by imposing the following condition: all roots of the AR and MA characteristic polynomials - $\Phi^{p}(\mathcal{B})$ and $\Theta^{q}(\mathcal{B})$, must lie outside the unit circle. Therefore an ARIMA(p,d,q)(p,d,q) process is stationary and invertible if the absolute value of all the roots of $\Phi^{p}(\mathcal{B})$ and $\Theta^{q}(\mathcal{B})$ is greater than one.
The stationarity constraint on the AR weights ensures that the modelled time series exhibits a stable and mean-reverting behaviour. It allows for the influence of past values to eventually decay with time, preventing the series from diverging to infinity. This can be demonstrated for a simple AR(1) model, for which the characteristic polynomial is
@@
$\epsilon_{t}=y_{t}-\theta_{1}y_{t-1}-\theta_{1}^{2}y_{t-2}-\theta_{1}^{3}y_{t-3}\ldots$ (15)
@@ 2.2 Nested Sampling
The Nested Sampling algorithm, first developed by Skilling (2006), is primarily used to compute the evidence term in Bayesian inference. Given a dataset $D$, and a model $\mathcal{M}$ characterized by parameters $\theta_{i}$, Bayesian inference updates the prior distribution on the model parameters $P(\theta_{i}|\mathcal{M})$ to obtain the posterior distribution $P(\theta_{i}|D;\mathcal{M})$ using Bayes’ theorem:
@@
It is generally intractable to analytically evaluate this integral, especially in higher-dimensional spaces. Nested Sampling is a numerical algorithm which transforms this multi-dimensional evidence integral into a one-dimensional problem by introducing the prior volume - the amount of fractional prior mass contained within an iso-likelihood contour of $\mathcal{L^{\prime}}$ in the prior space:
@@ 2.2.1 Sampler Output and Details
$\mathcal{D}_{\mathrm{KL}}$ is the Kullback-Leibler divergence (Kullback & Leibler, 1951) which represents the degree to which the posterior is concentrated relative to the prior:
@@
$n_{\text{live}}$ is the number of live points sampled from the prior. The total convergence time of the nested sampling run scales directly with $\mathcal{D}_{\mathrm{KL}}$ and $n_{\text{live}}$. For a fixed uncertainty $\sigma_{\log{Z}}$, $n_{\text{live}}$ is directly proportional to $\mathcal{D}_{\mathrm{KL}}$. The number of live points $n_{\text{live}}$ is therefore chosen to balance precision and time for convergence.
@@ 2.2.2 Model Selection
The evidence $Z_{i}$ associated with any model is thus a principled measure of the goodness of fit. This in turn is related to the likelihood function $\mathcal{L}(D|\theta_{i})$ as well as the size of the prior space over which the fit is performed. The latter factor ultimately acts as Occam’s razor that penalizes the evidence of an overfitting model.
@@ 3 Methodology
This section outlines the methodology adopted to apply the nested sampling algorithm for ARIMA models. We discuss details of the likelihood and prior functions in subsection 3.1, the nested sampler in subsection 3.2, and finally the model selection and validation procedures in subsections 3.3 and 3.4, respectively.
The following quantities are treated as the set of parameters characterizing any given ARIMA model: the respective AR and MA weights $\phi_{a}$ and $\theta_{m}$, the missing initial state ($D_{0}$ and $\epsilon_{0}$), the standard deviation $\sigma$ associated with the present error $\epsilon_{t}$ and the unconditional mean of the time series $\mu$, which can be shown to be related to the intercept term $c$ through:
@@ A Gaussian likelihood function is chosen for the $n$ observations of the given time series $D_{t}$ :
After fitting the ARIMA model, the resulting residuals are assumed to be independent Gaussian random variables with mean zero and constant variance ($\sigma^{2}$). Since the noise model enters explicitly through the likelihood function, the ARIMA forecast values $\mathbf{\hat{y}_{t}}$ used in computing the likelihood should be completely deterministic. These deterministic forecasts $\mathbf{\hat{y}_{t}}$ are obtained from Equation 3 by setting $\epsilon_{t}=0$. For simplicity, we treat the missing initial values for the time series data $D_{0}$ as nuisance parameters in our inference. The initial forecast errors $\epsilon_{0}$ are set to zero, thus evaluating a likelihood which is conditional on $\epsilon_{0}=0$.
@@
We now discuss the choice of prior distributions for the ARIMA model parameters. For $\sigma$, a reasonable choice of prior is a Half-Normal or a Half-Cauchy distribution with scale $\sigma^{\prime}$. We adopt a relatively weak prior for $\sigma$ by choosing a half-normal distribution truncated from below at zero and scaled by $\sigma^{\prime}=50$. The prior choice for the unconditional mean $\mu$ depends on the time series data under consideration. For a differenced or de-trended time-series, a standard normal distribution is chosen as prior for $\mu$. In other cases, a wide normal distribution centred around the expected long-term mean $\mu_{0}$ of the time-series is used. The prior distribution for the $p$ missing initial values of $D_{0}$ is set equivalent to the prior for $\mu$, as we expect these values to lie close to $\mu$ for a stationary series. For the initial residuals $\epsilon_{0}$, a standard normal distribution is chosen. In the case of an ARIMA($1,d,1$) model, a uniform distribution bounded between -1 and 1 is the most straightforward choice of priors for the coefficients $\phi_{a}$ and $\theta_{m}$. The bounds are in accordance with the stationarity and invertibility conditions discussed in subsection 2.1.1. However, for higher-order ARIMA models, the constraints on the individual weights are coupled and hence the choice of prior distributions for them is non-trivial.
@@
For this paper, we assign a mathematically constrained normal prior distribution $\mathcal{N}(0,\sigma^{\prime\prime})$ to the ARMA weights and implement a “rejection sampling" approach. The scale of the distribution $\sigma^{\prime\prime}$ is typically set to one as we do not expect the time-series under consideration to be very strongly auto-correlated. An initial set of $10^{3}\times n_{\text{live}}$ points is sampled from this distribution. The points are tested for stationarity and invertibility by calculating the roots of the AR and MA characteristic polynomials, and imposing the conditions discussed in subsection 2.1.1. The parameter points passing this test are then added to a pool of valid particles. This process is continued iteratively until the desired number of valid live particles $n_{\text{live}}$ is reached in this pool. The normal prior distributions on the weights are also constrained to the stationary and invertible regions of the parameter space.